suppose x = t^2 + t and y = t^3 + t^2. if x = 2 what are the two possible values of y.
Answer by htmentor(1343)

suppose x = t^2 + t and y = t^3 + t^2. if x = 2 what are the two possible values of y.
=================
If x = 2, we can solve for t:
t^2 + t = 2
t^2 + t - 2 = 0
Factor:
(t+2)(t-1) = 0
This gives t = -2, 1
So the possible values for y are:
(-2)^3 + (-2)^2 = -4
1 + 1 = 2
Ans: -4, 2